# Cooperative Cruising: Reinforcement Learning based Time-Headway Control for Increased Traffic Efficiency

### v1

@@ Abstract
The proliferation of Connected Automated Vehicles represents an unprecedented opportunity for improving driving efficiency and alleviating traffic congestion. However, existing research fails to address realistic multi-lane highway scenarios without assuming connectivity, perception, and control capabilities that are typically unavailable in current vehicles. This paper proposes a novel AI system that is the first to improve highway traffic efficiency compared with human-like traffic in realistic, simulated multi-lane scenarios, while relying on existing connectivity, perception, and control capabilities. At the core of our approach is a reinforcement learning based controller that dynamically communicates time-headways to automated vehicles near bottlenecks based on real-time traffic conditions. These desired time-headways are then used by Adaptive Cruise Control (ACC) systems to adjust their following distance. By (i) integrating existing traffic estimation technology and low-bandwidth vehicle-to-infrastructure connectivity, (ii) leveraging safety-certified ACC systems, and (iii) targeting localized bottleneck challenges that can be addressed independently in different locations, we propose a practical, safe, and scalable system that can positively impact numerous road users.
@@ 1 Introduction
Prior research on highway congestion reduction explored distributed and centralized vehicle speed control approaches. Distributed approaches typically implement an in-vehicle speed controller that uses information of the vehicle’s surroundings to decide when to increase headway and allow vehicles to merge into its lane. These approaches are scalable and effective when the location and time of lane-changes can be accurately predicted, such as in merge and certain bottleneck scenarios (Cui et al. 2021; Zhang et al. 2023a; Vinitsky et al. 2023). However, in multi-lane scenarios, lane changes can occur unpredictably at any time and any point on the road, driven by drivers’ intentions and behaviors, rendering these distributed approaches ineffective. Centralized approaches leverage aggregate traffic data to provide high-level guidance for influencing spatio-temporal traffic density (Bayen et al. 2020). While this circumvents the need to predict lane changes, these approaches face challenges in generating the complex speed-control commands necessary to achieve the desired density patterns.
To address these challenges, we propose a centralized AI system that influences density more directly by generating time-headway requests that are used by ACC systems to adjust their vehicles’ headways. At the core of our systems is a reinforcement learning based controller that continually outputs desired vehicle headways for each road segment leading to the bottleneck, based on real-time traffic conditions. To narrow the gap to real-world deployment, our system is designed to integrate with existing traffic estimation technology, low-bandwidth vehicle-to-infrastructure connectivity, and safety-certified ACC systems. As shown in hundreds of large-scale experiments involving hundreds of vehicles and various CAV penetration rates, our system is the first to demonstrate significant traffic flow improvements over human-driven traffic in realistic, simulated multilane scenarios, where previous methods fall short. As a secondary contribution to improving traffic flow evaluation accuracy, we identify a flaw in a commonly used throughput metric and introduce a modified average speed metric that accounts for entry delays caused by congestion, addressing a limitation highlighted in prior work (Cui et al. 2021). By leveraging existing technologies, ensuring safety through ACC systems, and addressing localized bottleneck challenges, we propose a practical, safe, and scalable system that can potentially enhance the travel experience of numerous road users. Our code is publicly available on GitHub.
@@ 2 Related Work
Traffic congestion poses a significant challenge in highway planning, prompting the development of traffic flow models to understand and mitigate its adverse impacts (Hall 1992; Ni and Leonard 2006; van Wageningen-Kessels et al. 2015; Ferrara, Sicona, and Silvia 2018; Mohammadian et al. 2021). Studies have shown that traffic breakdowns can occur spontaneously and stochastically, even without the presence of bottlenecks (Sugiyama et al. 2008), a phenomenon that classical traffic flow theories fail to adequately explain (Kerner et al. 2015; Kerner 2016). Traffic microsimulators, such as SUMO (Krajzewicz et al. 2012) which we use in this paper, were thus suggested for reproducing observable traffic phenomena, using car-following models of human driving. One such model is the Intelligent Driver Model (IDM), which has been instrumental in reproducing complex phenomena like traffic jams, stop-and-go waves, and bottleneck congestion (Treiber, Hennecke, and Helbing 2000). IDM can reproduce realistic vehicle interactions, particularly when modeling diverse traffic scenarios, and is superior to other models such as the Optimal Velocity Model (OVM) (Bandō et al. 1994, 1995) that tends to smooth out traffic by encouraging vehicles to maintain an optimal velocity.
@@
The idea to utilize automated vehicles as mobile actuators for alleviating traffic problems is based on the assumption that these vehicles can be systematically coordinated and controlled to optimize traffic flow (Stern et al. 2018; Wu, Bayen, and Mehta 2018; Delle Monache et al. 2019; Gora et al. 2020; Wang et al. 2022, 2023). Inspired by this idea, the CIRCLES project (CIRCLES 2020) has focused on developing traffic control algorithms to optimize traffic flow on highways, and conducted a large-scale, open-road field experiment with 100 CAVs (Wang et al. 2024; Lee et al. 2024).
@@
Since developing model-based controllers for traffic congestion is challenging due to the problem’s scale and complexity, research on data-driven controllers trained in simulation using Reinforcement Learning (RL) (Sutton and Barto 2018) has emerged. This line of research demonstrated that centralized RL controllers enable wave dissipation and significant average speed increase in single-lane roads with merges or bottlenecks with as low as 10% CAV penetration (Kreidieh, Wu, and Bayen 2018; Vinitsky et al. 2018a, b, c). However, deployment of these approaches might not be scalable due to the challenge of learning a policy that individually controls a large numbers of vehicles.
@@
To enable scalability, distributed approaches have employed in-vehicle speed controllers that utilize information about a vehicle’s surroundings to determine when to increase headway and allow other vehicles to merge into its lane. These approaches work well when the location and time of lane-changes can be accurately predicted, e.g. in merge and certain bottleneck scenarios (Cui et al. 2021; Zhang et al. 2023a; Vinitsky et al. 2023). However, in multi-lane scenarios, lane changes can occur unpredictably at any time and point on the road, rendering these distributed approaches ineffective.
@@
As an alternative, centralized approaches have used aggregate traffic data to provide high-level guidance without modeling local behaviors, such as lane-changes. The Variable Speed Limit (VSL) method (Hegyi, De Schutter, and Hellendoorn 2005; Lu and Shladover 2014) aims to prevent traffic breakdowns by regulating inflow into congestion-prone areas. Extensive studies of VSL (Li et al. 2017; Nie et al. 2021; Alasiri, Zhang, and Ioannou 2023; Zhang et al. 2023b; Hua and Fan 2023) have explored its application through optimization problems, such as minimizing total travel time or maximizing traffic throughput (Bayen et al. 2022). However, these approaches have struggled to address realistic multi-lane scenarios due to the difficulty of achieving the required road dynamics when only speed limits are controlled (see Section 4.2). To overcome these limitations, we propose a centralized AI system that avoids the need to predict lane changes or implement complex speed control. Instead, our method employs time-headway control (Pang and Huang 2022) to directly influence spatio-temporal traffic density to enhance traffic efficiency. By leveraging existing traffic estimation technology, low-bandwidth vehicle-to-infrastructure communication, and safety-certified ACC systems, our approach offers a practical, safe, and scalable solution for real-world applications.
@@ 3 Domain Description
In this section, we define the problem addressed by this paper, and the simulation setup used in the experiments.
@@ 3.1 Problem Description
Our problem is defined as follows. Given a road network with multiple lanes, a merging road, and mixed autonomy traffic consisting of both human-driven vehicles and CAVs, maximize the network’s traffic efficiency by controlling CAVs, where traffic efficiency is measured in terms of average speed. We assume that CAVs are altruistic, sharing the common goal of reducing traffic congestion, which can be facilitated by incentivizing such behavior. A solution to our problem is a control policy that maps the traffic state to actions that influence CAVs to enhance traffic efficiency. For reasons described in Section 4, we propose and focus on control policies that influence CAVs equipped with technologies such as ACC, by sending them time-headway commands based on real-traffic information.
We note that in simulated open road networks where vehicles enter and exit dynamically, increased average speed may not accurately reflect real-world traffic flow improvements due to discrepancies between finite simulated roads and real-world conditions (Cui et al. 2021). To address this, we propose a novel average speed metric that aligns more closely with real-world scenarios, enabling more reliable evaluation of our system’s performance (Section 4.1).
@@ 3.2 Simulation Platform & Scenario Parameters
The road network used in our simulations is a 2 km segment of I-24 in Tennessee, USA, extracted from OpenStreetMap (2017). The simulation includes hundreds of vehicles engaging in complex lane-changing and car-following interactions, with vehicles merging onto the highway and executing required lane changes. Figure 1a presents a partial snapshot of our scenario in SUMO, featuring a centralized control policy that senses traffic conditions and issues time-headway commands. These commands are utilized by ACC systems to dynamically adjust vehicle headways.
Human-driven vehicles are modeled using IDM (Treiber, Hennecke, and Helbing 2000) which has a fixed time-headway parameter and a safety enforcement module. CAVs are modelled using IDM with a time-headway parameter that can be dynamically adjusted. Before real-world deployment, automated vehicle models should be calibrated with real-world data prior to using them with time-headway controllers, however due to the lack of such data and the aim of ACC systems to follow human-driving behaviors, we used the aforementioned adjustable IDM models. An additional challenge is modeling lane-change behavior, which significantly affects traffic flow (Figure 1b). Aggressive lane-change behavior, characterized by vehicles merging into smaller gaps, causes major disturbances and decreases both speed and throughput. Timid lane-change behavior cause disturbances by low utilization of additional lanes. We adjusted lane-change aggressiveness through visual inspection to better align with real-world driving behavior. Lane-change behavior should ideally also be calibrated with highway data if available. Values of all SUMO parameters can be found in the Appendix.
@@ 4 Methodology
In this section, we present our methodology for designing and testing the proposed time-headway controllers. First, we introduce a congestion metric that overcomes the limitations of finite-length simulations highlighted in prior research. Next, we analyze vehicle interactions contributing to traffic inefficiencies, identify the shortcomings of existing methods, and justify our centralized time-headway control approach. We then outline our practical design choices. Finally, we describe a baseline fixed-value time-headway control policy, and our proposed RL-based control policy.
@@ 4.1 Congestion Metrics
To demonstrate improvements in traffic flow efficiency, it is essential to define how congestion is measured. In an ideal scenario with constant inflows, there are multiple metrics that would all lead to the same quality-ordering of congestion-reduction control policies: maximizing average speed, maximizing network outflow, and minimizing average time delay (Dresner 2008). Both average speed and average throughput are commonly used metrics, but each has its drawbacks in simulations. Average throughput is sensitive to the simulation length, since a temporary decrease in throughput can be offset by a subsequent increase, if the simulation is long enough and includes periods with demand lower than road capacity. While changes in average speed can effectively measure congestion, caution is needed when using this metric in finite road length simulations, especially when testing controllers. A controller that prevents vehicles from entering the simulation might artificially inflate speeds without being penalized for the reduction of speeds that would have happened in real-world road sections preceding the simulated road network  (Cui et al. 2021).
We solve the aforementioned problem by tracking vehicles whose entry to the simulation is delayed and penalizing for the delay time. The metric we use is the average speed change compared to a simulated human-driven traffic, taking into account the delay time of each of the vehicles:
$\Delta V\equiv\frac{1}{N}\sum_{i}^{N}\frac{\bar{v}_{\text{control}}^{\left(i\right)}-\bar{v}_{\text{baseline}}^{\left(i\right)}}{\bar{v}_{\text{baseline}}^{\left(i\right)}}$ (1)
where $N$ is the total number of vehicles, $\bar{v}_{\text{baseline}}^{\left(i\right)}$ is vehicle $i$’s average speed in simulated human-driven traffic, and $\bar{v}_{\text{control}}^{\left(i\right)}$ is the same vehicle’s speed in the controlled simulation.
@@
$\bar{v}^{\left(i\right)}=\frac{L^{\left(i\right)}}{\min\left(T_{f}^{\left(i\right)},T_{sim}\right)-T_{s}^{\left(i\right)}}$ (2)
where $L^{\left(i\right)}$ is the distance driven by vehicle $i$, $T_{s}^{\left(i\right)}$ is its planned simulation entry time, $T_{f}^{\left(i\right)}$ is its measured finish time, and $T_{sim}$ is the final simulation time. $T_{s}^{\left(i\right)}$ of each vehicle is known apriori and stored in a per-vehicle route file used to initialize the SUMO simulation.
@@ 4.2 Centralized Time-Headway Control
To motivate our centralized time-headway control approach, we identify and address factors that may have prevented prior methods from handling realistic multi-lane scenarios. Prior research has shown that lane changes during merges can cause slowdowns and trigger congestion. However, if a vehicle in the target lane preemptively increases its headway as it approaches the merge, it can prevent congestion from forming, as illustrated in Figure 2a. Distributed approaches typically implement an in-vehicle speed controller that uses information of the vehicle’s surroundings to decide when to preemptively increase headway. These approaches are scalable and effective when the location and time of lane-changes can be accurately predicted, such as in merge and certain bottleneck scenarios. However, in multi-lane scenarios, lane changes can occur unpredictably at any time and any point on the road, driven by drivers’ intentions and behaviors, rendering these distributed approaches ineffective.
In such scenarios, it may be desirable to try approaches of higher-level control over road density, motivated by real-world data showing that lower traffic density reduces the negative impact of lane changes on traffic flow (Yang, Wang, and Quddus 2019; Gao and Levinson 2023). Since density is an aggregate measure, it is naturally controlled using a centralized controller. Centralized approaches that attempt to control each CAV individually face scalability issues (Cui et al. 2021). Instead, centralized approaches such as Variable Speed Limit (VSL) used aggregate traffic data to provide high-level guidance by employing speed-limit adjustments in controlled road segments. However, these approaches struggle with controlling traffic density. As shown in the qualitative vehicle trajectory time-space diagram in Figure 2b, applying a constant speed limit in a road segment can only temporarily reduce downstream density (blue area). In contrast, constant time-headway signals maintain lower downstream density for an arbitrary duration. This is empirically demonstrated in Figure 2c, where a constant time-headway control signal results in a complex, time-dependent upstream speed profile that would be difficult to achieve with speed-limit control.
Therefore, we develop a centralized time-headway control capable of reducing traffic density around bottlenecks, allowing vehicles to self-organize in the sparser traffic, as often observed in real-world uncontrolled traffic. Our system uses high-level traffic information to control vehicles near road merges, enabling scalable and independent deployment at numerous highway junctions.

### v2

# Cooperative Cruising: Reinforcement Learning-Based Time-Headway Control for Increased Traffic Efficiency

@@ Abstract
The proliferation of connected automated vehicles represents an unprecedented opportunity for improving driving efficiency and alleviating traffic congestion. However, existing research fails to address realistic multi-lane highway scenarios without assuming connectivity, perception, and control capabilities that are typically unavailable in current vehicles. This paper proposes a novel AI system that is the first to improve highway traffic efficiency compared with human-like traffic in realistic, simulated multi-lane scenarios, while relying on existing connectivity, perception, and control capabilities. At the core of our approach is a reinforcement learning based controller that dynamically communicates time-headways to automated vehicles near bottlenecks based on real-time traffic conditions. These desired time-headways are then used by adaptive cruise control (ACC) systems to adjust their following distance. By (i) integrating existing traffic estimation technology and low-bandwidth vehicle-to-infrastructure connectivity, (ii) leveraging safety-certified ACC systems, and (iii) targeting localized bottleneck challenges that can be addressed independently in different locations, we propose a potentially practical, safe, and scalable system that can positively impact numerous road users.
@@ 1 Introduction
Prior research on highway congestion reduction explored distributed and centralized vehicle speed control approaches. Distributed approaches typically implement an in-vehicle speed controller that uses information of the vehicle’s surroundings to decide when to increase headway and allow vehicles to merge into its lane. These approaches are scalable and effective when the location and time of lane-changes can be accurately predicted, such as in merge and certain bottleneck scenarios (Cui et al. 2021; Zhang et al. 2023a; Vinitsky et al. 2023). However, in multi-lane scenarios, lane changes can occur unpredictably at any time and any point on the road, driven by drivers’ intentions and behaviors, rendering these distributed approaches ineffective. Centralized approaches leverage aggregate traffic data to provide high-level guidance for influencing spatio-temporal traffic density (Bayen et al. 2020). While such approaches circumvent the need to precisely predict lane changes, they face challenges in generating the complex speed-control commands needed for influencing traffic density to reduce congestion.
To address these challenges, we propose a centralized AI system that influences density more directly by generating time-headway requests that are used by ACC systems to adjust their vehicles’ headways. At the core of our systems is a reinforcement learning based controller that continually outputs desired vehicle headways for each road segment leading to the bottleneck, based on real-time traffic conditions. To narrow the gap to real-world deployment, our system is designed to integrate with existing traffic estimation technology, low-bandwidth vehicle-to-infrastructure connectivity, and safety-certified ACC systems. Through hundreds of large-scale simulated experiments, we show that our system significantly improves traffic flow compared to human-driven traffic, particularly in scenarios where previous methods fall short. As a secondary contribution, we address a previously identified flaw in measuring average speed in simulated road networks where vehicles enter and leave the simulation dynamically (Cui et al. 2021), and introduce a method to better approximate realistic average speed values. By leveraging existing technologies, ensuring safety through ACC systems, and addressing localized bottleneck challenges, we propose a potentially practical, safe, and scalable system that could enhance the travel experience of numerous road users. Our code is publicly available on GitHub [^1] https://coopcruise.github.io/.
@@ 2 Related Work
Traffic congestion poses a significant challenge in highway planning, prompting the development of traffic flow models to understand and mitigate its adverse impacts (Hall 1996; Ni and Leonard II 2006; Ferrara et al. 2018; Mohammadian et al. 2021). Studies have shown that traffic breakdowns can occur spontaneously and stochastically, even without the presence of bottlenecks (Sugiyama et al. 2008), a phenomenon that classical traffic flow theories fail to adequately explain (Kerner 2016). Traffic microsimulators, such as SUMO (Krajzewicz et al. 2012) which we use in this paper, were suggested for reproducing observable traffic phenomena, using car-following models of human driving. One such model is the Intelligent Driver Model (IDM), which has been instrumental in reproducing complex phenomena like traffic jams, stop-and-go waves, and bottleneck congestion (Treiber, Hennecke, and Helbing 2000). IDM can reproduce realistic vehicle interactions, particularly when modeling diverse traffic scenarios, and is superior to other models such as the Optimal Velocity Model (Bandō et al. 1995) that tends to smooth out traffic by encouraging vehicles to maintain optimal velocity.
@@
The idea to utilize automated vehicles as mobile actuators for alleviating traffic problems is based on the assumption that these vehicles can be systematically coordinated and controlled to optimize traffic flow (Stern et al. 2018; Wu, Bayen, and Mehta 2018; Delle Monache et al. 2019; Wang et al. 2023a, b). Inspired by this idea, the CIRCLES project (CIRCLES 2020) has focused on developing traffic control algorithms to optimize traffic flow on highways, and conducted a large-scale, open-road field experiment with 100 CAVs (Wang et al. 2024a; Lee et al. 2024).
@@
Since developing model-based controllers for traffic congestion is challenging due to the problem’s scale and complexity, research on data-driven controllers trained in simulation using reinforcement learning (RL) (Sutton and Barto 2018) has emerged. This line of research demonstrated that centralized RL controllers enable wave dissipation and significant average speed increase in single-lane roads with merges or bottlenecks with as low as 10% CAV penetration (Kreidieh, Wu, and Bayen 2018; Vinitsky et al. 2018; Wang et al. 2024b). However, these approaches may not scale well, as learning a single policy to individually control numerous vehicles becomes infeasible due to the high-dimensional action space.
@@
To enable scalability, distributed approaches have employed in-vehicle speed controllers that utilize information about a vehicle’s surroundings to determine when to increase headway and allow other vehicles to merge into its lane. These approaches work well when the location and time of lane-changes can be accurately predicted, e.g. in merge and certain bottleneck scenarios (Cui et al. 2021; Zhang et al. 2023a; Vinitsky et al. 2023). However, these approaches become ineffective in multi-lane scenarios, where lane changes occur unpredictably at any time and place.
@@
As an alternative, centralized approaches have used aggregate traffic data to provide high-level guidance without modeling local behaviors such as lane-changes. The Variable Speed Limits (VSL) method (Hegyi, De Schutter, and Hellendoorn 2005; Lu and Shladover 2014) aims to prevent traffic breakdowns by regulating inflow into congestion-prone areas. Extensive studies of VSL (Alasiri, Zhang, and Ioannou 2023; Zhang et al. 2023b; Hua and Fan 2023) have explored its application through optimization problems, such as minimizing total travel time or maximizing traffic throughput (Bayen et al. 2022). However, these approaches have struggled to address realistic multi-lane scenarios due to the difficulty of achieving the required road dynamics when only speed limits are controlled (as discussed in Section 4.2). To overcome these limitations, we propose a centralized AI system that avoids the need to predict lane changes or implement complex speed control. Instead, our method employs time-headway control to directly influence spatio-temporal traffic density to enhance traffic efficiency. Prior work (Pang and Huang 2022) proposed using time headways for platoon leaders, whereas our method assigns them to road segments.
@@ 3 Domain Description
In this section, we define the problem addressed by this paper and the simulation setup used in the experiments.
@@ 3.1 Problem Description
Our problem is defined as follows. Given a road network with multiple lanes, a merging road, and mixed autonomy traffic consisting of both human-driven vehicles and CAVs, maximize the network’s traffic efficiency by controlling CAVs, where traffic efficiency is measured in terms of average speed. We assume that CAVs are altruistic, sharing the common goal of reducing traffic congestion, which can be facilitated by incentivizing such behavior. A solution to our problem is a control policy that maps the traffic state to actions that influence CAVs to enhance traffic efficiency. For reasons described in Section 4, we propose and focus on control policies that influence CAVs equipped with technologies such as ACC, by sending them time-headway commands based on real-time traffic information.
We note that in simulated open road networks where vehicles enter and exit dynamically, increased average speed may not reflect real-world traffic flow improvements due to discrepancies between finite simulated roads and real-world conditions (Cui et al. 2021). We therefore propose a novel average speed metric that aligns more closely with real-world scenarios, enabling more reliable evaluations of average speed improvements in open road simulations (Section 4.1).
@@ 3.2 Simulation Platform & Scenario Parameters
The road network used in our simulations is a 2 km segment of I-24 in Tennessee, USA, extracted from OpenStreetMap (2017), chosen as a representative example of a general highway merge geometry. The simulation includes hundreds of vehicles engaging in complex lane-changing and car-following interactions, with vehicles merging onto the highway and executing required lane changes. Prior studies (Samaei et al. 2023) modeled a larger section of I-24 using real-world traffic data. Figure 1a presents a partial snapshot of our scenario in SUMO, featuring a centralized control policy that senses traffic conditions and issues time-headway commands. These commands are utilized by ACC systems to dynamically adjust vehicle headways.
Human-driven vehicles are modeled using IDM (Treiber, Hennecke, and Helbing 2000) which has a constant time-headway parameter and a safety enforcement module. CAVs are modelled using IDM with a time-headway parameter that can be dynamically adjusted. Before real-world deployment, automated vehicle models should be calibrated with real-world data prior to using them with time-headway controllers, however due to the lack of such data and the aim of ACC systems to follow human-driving behaviors, we used the aforementioned adjustable IDM models. An additional challenge is modeling lane-change behavior, which significantly affects traffic flow (Figure 1b). Aggressive lane-change behavior, characterized by vehicles merging into smaller gaps, causes major disturbances and decreases both speed and throughput. Timid lane-change behavior cause disturbances by low utilization of additional lanes. We adjusted lane-change aggressiveness through visual inspection to better align with real-world driving behavior. Lane-change behavior should ideally also be calibrated with highway data if available. Values of all SUMO parameters can be found in the Appendix.
@@ 4 Methodology
In this section, we present our methodology for designing and testing the proposed time-headway controllers. First, we introduce an average-speed metric that addresses the limitations of average-speed measurements in simulated open road networks which were pointed out by prior research. Next, we analyze vehicle interactions contributing to traffic inefficiencies, identify the shortcomings of existing methods, and justify our centralized time-headway control approach. We continue by outlining our practical design choices. Finally, we describe a baseline fixed-value time-headway control policy, and our proposed RL-based control policy.
@@ 4.1 Congestion Metrics
To demonstrate improvements in traffic flow efficiency, it is essential to define how congestion is measured. Average speed, average throughput, and time delay with respect to travel time at maximum speed are commonly used metrics, but have drawbacks in simulations. Average throughput is sensitive to simulation length, since a temporary throughput decrease can be offset by a subsequent increase, for long enough simulations including periods with demand lower than road capacity. While changes in average speed or time delay can effectively measure congestion, caution is needed when using these metrics in finite road length simulations: controllers which prevent vehicles from entering the simulation might artificially inflate speeds without being penalized for the reduction of speeds that would have happened in real-world road sections preceding the simulated road network  (Cui et al. 2021).
We solve the aforementioned problem by tracking vehicles whose entry to the simulation is delayed and penalizing for their delay time. The metric we use is the average speed change compared to a simulated human-driven traffic, taking into account the entry delay time of each of the vehicles:
$\Delta V\equiv\frac{1}{N}\sum_{i}^{N}\frac{\bar{v}_{\text{control}}^{\left(i\right)}-\bar{v}_{\text{baseline}}^{\left(i\right)}}{\bar{v}_{\text{baseline}}^{\left(i\right)}},$ (1)
where $N$ is the total number of vehicles, $\bar{v}_{\text{baseline}}^{\left(i\right)}$ is vehicle $i$’s average speed in simulated human-driven traffic, and $\bar{v}_{\text{control}}^{\left(i\right)}$ is the same vehicle’s speed in a simulation with headway control. Note that this metric takes all vehicles into account, ensuring that traffic-blocking strategy is not a viable solution.
@@
$\bar{v}^{\left(i\right)}=\frac{L^{\left(i\right)}}{\min\left(T_{f}^{\left(i\right)},T_{sim}\right)-T_{s}^{\left(i\right)}},$ (2)
where $L^{\left(i\right)}$ is the distance driven by vehicle $i$, $T_{s}^{\left(i\right)}$ is its planned simulation entry time, $T_{f}^{\left(i\right)}$ is its measured exit time, and $T_{sim}$ is the final simulation time. $T_{s}^{\left(i\right)}$ of each vehicle is known apriori and stored in a per-vehicle route file used to initialize the SUMO simulation.
@@ 4.2 Centralized Time-Headway Control
To motivate our centralized time-headway control approach, we identify and address factors that may have prevented prior methods from handling realistic multi-lane scenarios. Prior research has shown that lane changes during merges can cause slowdowns and trigger congestion. However, if a vehicle in the target lane preemptively increases its headway as it approaches the merge, it can prevent congestion from forming, as illustrated in Figure 2a. Distributed approaches typically implement an in-vehicle speed controller that uses information of the vehicle’s surroundings to decide when to preemptively increase headway. These approaches are scalable and effective when the location and time of lane-changes can be accurately predicted, such as in merge and certain bottleneck scenarios. However, in multi-lane scenarios, lane changes can occur unpredictably at any time and place on the road, driven by drivers’ intentions and behaviors, rendering these distributed approaches ineffective.
In such scenarios, it may be desirable to try approaches for higher-level control over road density, motivated by real-world data showing that lower traffic density reduces the negative impact of lane changes on traffic flow (Yang, Wang, and Quddus 2019; Gao and Levinson 2023). Since density is an aggregate measure, it is naturally controlled using a centralized controller. Centralized approaches that attempt to control each CAV individually face scalability issues (Cui et al. 2021). Instead, centralized approaches such as Variable Speed Limits used aggregate traffic data to provide high-level guidance by employing speed-limit adjustments in controlled road segments. However, these approaches struggle with controlling traffic density. The qualitative vehicle trajectory time-space diagram in Figure 2b shows that applying a constant speed limit in a road segment can reduce downstream density only temporarily (blue area). In contrast, constant time-headway signals maintain lower downstream density for an arbitrary duration. This is empirically demonstrated in Figure 2c, where a constant time-headway control signal results in a complex, time-dependent upstream speed profile that would be difficult to achieve with speed-limit control.
We develop a centralized time-headway control capable of reducing traffic density around bottlenecks, allowing vehicles to self-organize in sparser traffic, as often observed in real-world uncontrolled traffic. Our system offers a scalable solution, as its use of local traffic information and control enables independent deployment at numerous highway junctions.